OpenStudy (linn99123):

NEED HELP FAST GOT 4 MINUTES LEFT WILL FAN AND GIVE MEDAL :o A muffin recipe calls for 4 tablespoons of water to make 10 servings. How many tablespoons of water are needed to make 8 servings? A. 2.5 tablespoons B. 3.2 tablespoons C. 5 tablespoons D. 20 tablespoons

OpenStudy (anonymous):

c

OpenStudy (anonymous):

I'd say 3.2 just based on deduction skills.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

4/10= .4 .4*8=3.2

OpenStudy (anonymous):

Any more questions?

OpenStudy (anonymous):

The Answer Is B! 4 / 10 = .4 . 4 * 8 = 3 . 2

OpenStudy (anonymous):

I have one!! can you help me @OrangeMaster? http://openstudy.com/updates/52e82a5ee4b042661d4c904f